OpenStudy (anonymous):

Been stuck for hours The function H(t) = −16t2 + 90t + 50 shows the height H(t), in feet, of a projectile after t seconds. A second object moves in the air along a path represented by g(t) = 28 + 48.8t, where g(t) is the height, in feet, of the object from the ground at time t seconds. Part A: Create a table using integers 1 through 4 for the 2 functions. Between what 2 seconds is the solution to H(t) = g(t) located? How do you know? (6 points) Part B: Explain what the solution from Part A means in the context of the problem. (4 points)

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Anybody got an idea? I'm basically just needing help on part A.

OpenStudy (mathstudent55):

Start by evaluating both functions at t = 1, t = 2, t = 3, t = 4

OpenStudy (anonymous):

16 90 28 48?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

How did you get 16?

OpenStudy (mathstudent55):

H(t) = −16t2 + 90t + 50 Let t = 1. What do you get for H(1) ?

OpenStudy (anonymous):

I just found it, idk how i got it, I guessed on it

OpenStudy (anonymous):

How do i find h(1)

OpenStudy (anonymous):

Like, i'm 100% lost on how I make these, I found some functions but idk how to graph them, can you explain?

OpenStudy (mathstudent55):

H(t) = −16t2 + 90t + 50 To find H(1) you replace t with 1, and you evaluate the expression. H(1) = -16(1)^2 + 90(1) + 50

OpenStudy (mathstudent55):

H(1) = -16 + 90 + 50 = 124 Now you need H(2). Replace t with 2 in the H(t) function and evaluate the right side. H(2) = -16(2^2) + 90(2) + 50

OpenStudy (anonymous):

How do I evaluate it? Like how'd you get that

OpenStudy (anonymous):

1 is H(1) = -16 + 90 + 50 = 124, I need to solve for 2,3, and 4?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

I need to find 2,3,4 but how?

OpenStudy (mathstudent55):

Evaluating an expression is just using arithmetic to calculate what it is equal to.

OpenStudy (mathstudent55):

For t = 2, you get H(2) = -16(2^2) + 90(2) + 50 What is the right side equal to?

OpenStudy (anonymous):

H(2) = -16(2v2) + 180 + 50?

OpenStudy (mathstudent55):

If I asked you what is 2 + 2, would you answer 2 + 2, or would you answer 4?

OpenStudy (anonymous):

H(2) -16^4 + 230?

OpenStudy (mathstudent55):

good, keep going

OpenStudy (anonymous):

H(2) - 64 + 230?

OpenStudy (mathstudent55):

good

OpenStudy (mathstudent55):

H(2) = - 64 + 230?

OpenStudy (anonymous):

So it'd be H(2) = 294, or since 64 is negative would it be H(2) = 166?

OpenStudy (mathstudent55):

H(2) = 166

OpenStudy (mathstudent55):

So far we have: H(1) = 124 H(2) = 166 Now we still need H(3) and H(4) for function H(t) H(t) = −16t2 + 90t + 50 H(3) = -16(3)^2 + 90(3) + 50

OpenStudy (anonymous):

H(3) = -96 + 320

OpenStudy (anonymous):

H(3) = 224?

OpenStudy (mathstudent55):

No. First do 3^2

OpenStudy (anonymous):

9

OpenStudy (mathstudent55):

H(3) = -16(3)^2 + 90(3) + 50 H(3) = -16(9) + 90(3) + 50 Now do -16 * 9. You already have 90(3) + 50 = 320

OpenStudy (anonymous):

-144

OpenStudy (anonymous):

-144 + 270 + 50 right?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

H(3)=176?

OpenStudy (mathstudent55):

Correct. Now we need H(4) H(4) = -16(4)^2 + 90(4) + 50 Remember to start with 4^2 then multiply that by -16

OpenStudy (anonymous):

156?

OpenStudy (mathstudent55):

I got 154 H(4) = -16(4)^2 + 90(4) + 50 = -256 + 360 + 50 =

OpenStudy (mathstudent55):

Ok. We have now function H(t) evaluated at t = 1, t = 2, t = 3, t = 4. Now we need to do the same for function g(t)

OpenStudy (anonymous):

Four more for g(t)?

OpenStudy (anonymous):

Alright, what's the functions?

OpenStudy (anonymous):

You there math?

OpenStudy (mathstudent55):

The g(t) function is simpler. Just multiply the number (1, 2, 3, 4) by 48.8, then add to 28. g(t) = 28 + 48.8t g(1) = 28 + 48.8(1) = g(1) = 28 + 48.8(2) = g(1) = 28 + 48.8(3) = g(1) = 28 + 48.8(4) =

OpenStudy (anonymous):

g(1) =76.8 g(2)=125.6 g(3)= 174.4 g(4)= 223.2

OpenStudy (mathstudent55):

Correct. Great job on function g(t). Now we place all values in a table like part A asks for.